OpenStudy (anonymous):

Help me Please! I will give medals. Write an equation that can be used to solve: Emilio has 10 more nickels than dimes, and the total value of his coins is $2.90. How many nickels does he have? Let n represent the number of nickels.

OpenStudy (anonymous):

i think the answer would be he has 10 nickels and four dimes

OpenStudy (anonymous):

Let n represent the number of nickels. Let d represent the number of dimes. n = d+10 5n + 10d = 290 cents

OpenStudy (anonymous):

1 nickle is 5 cents and 4 nickles is 20 cents.

OpenStudy (anonymous):

sorry i was thinking of a quarter

OpenStudy (anonymous):

It's ok :)

OpenStudy (anonymous):

do you know how many dimes he has

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@sangya21 I'm still not quite sure on how to work that out, I'm kind of confused...

OpenStudy (anonymous):

40 nickels= $2.00 and then add 6 more nickels which will get $2.30 then 6 dimes which will give you $2.90

OpenStudy (anonymous):

does that make scene

OpenStudy (anonymous):

n = d+10 5n + 10d = 290 cents use n = d+10 in second eq. to find d then use that value of d in first eq to find n.

OpenStudy (anonymous):

@DawnStar53 does my answer make scene

OpenStudy (anonymous):

Would it be 25? Because the value of a nickle is 5 and you multiply that by 5? 25+10d=290?

OpenStudy (anonymous):

who are you talking to @DawnStar53

OpenStudy (anonymous):

5(d+10) + 10d = 290 5d + 50 + 10d = 290 15d = 290 - 50 15 d = 240 d = 240/15 = 16 total number of dimes = 16 as number of nickels is 10 more than number of dimes then, total number of nickel = 16+ 10 = 26

OpenStudy (anonymous):

It makes a little bit of sense but it still confuses me.

OpenStudy (anonymous):

never mind i guess i am just confusing myself

OpenStudy (anonymous):

@DawnStar53 did you get it?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Oh, nevermind :) Thank you @sangya21

OpenStudy (anonymous):

where are you getting confused?

OpenStudy (anonymous):

I didn't understand the distributing, but I understand it now! :)

OpenStudy (anonymous):

Ok :)